These were accepted.
Will put this data on Satabeast.
Expected to meet target of end of May.
Deferred for 6 months.
To be removed from list, blog article tracked elsewhere.
Neil has done more work on this and will be sending emails to users/groups shortly.
List updated and emails sent to users.
Alastair has emailed Tony but no reply.
A list of printers has been sent. It all seems to work so now waiting response from IS.
Still not complete. Not mandatory, but pass details to Carol if you haven't done so.
Blog articles discussed:
Blog articles suggested:
Report from Computing Executive Group
7 servers have been ordered. 5 have arrived subsequent to the meeting.
Reports from units.
It was noted that there might be spare memory available for heffalump from an R900 that is about to be retired.
It was agreed that it would be useful to provide a list of public servers. US will arrange.
It was noted that the MFD replacements will be the responsibility of the Services Unit.
We also discussed the need for destroying disks when recycling/re-allocating. For now, we agreed that, in addition to all server disks, we would also wipe all disks from admin machines. US will write a procedure. Stephen will check which wires dban is availabel on - we think it is all wires.
Topics for discussion
It is important that Units complete the trawl through Neil's list and identify sites that they are responsible for. Users have now been contacted by email regarding individual sites. Neil explained what he now plans to do with regards to user homepages and group files. Neil will write a blog article and users will be contacted individually. As further advice appears, this will be added to the wiki page.
The next meeting will be on Wednesday 30th May 2012 in 4.31. George will convene.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh